# Improved Constructions and Lower Bounds for Maximally Recoverable Grid Codes

### v1

@@ IV-A $n^{\lceil h/2\rceil}$ Lower Bound for the Gabidulin Construction
For the lower bound, we assume without loss of generality that $m=2$. Label the two left vertices as $\alpha$ and $\beta$ and label the right vertices as $1,\ldots,n$. If $h$ is odd then we claim that for any $(h+1)/2$ vertices $j_{1},\ldots,j_{(h+1)/2}\in[n]$ the sum of $\gamma_{\alpha,j_{1}}-\gamma_{\beta,j_{1}}\ldots+\gamma_{\alpha,j_{(h+1)/2}}-\gamma_{\beta,j_{(h+1)/2}}$ is distinct for distinct choices of right vertices. We note this claim would then force the field size to be at least $\binom{n}{(h+1)/2}$.
@@
$\gamma_{\alpha,i_{1}}-\gamma_{\beta,i_{1}}+\ldots+\gamma_{\alpha,i_{k}}-\gamma_{\beta,i_{k}}-(\gamma_{\alpha,j_{1}}-\gamma_{\beta,j_{1}}+\ldots+\gamma_{\alpha,j_{k}}-\gamma_{\beta,j_{k}})=0$ (4)
for $k\leq(h+1)/2$. Consider the following spanning tree with $i_{1}$ as a root: the node $i_{1}$ is connected to both left vertices and all the other right nodes are connected to the same left node say $\alpha$. If we consider the edges $(x,\beta)$ for $x=i_{2},\ldots,i_{k},j_{1},\ldots,j_{k}$, the four cycle involving it is $C_{j_{\ell}}$ is $j_{\ell},\beta,j_{1},\alpha$. $\gamma_{C_{j_{\ell}}}$ for $\ell=2,\ldots,h+1$ must be linearly independent and in particular their sum must be non-zero. We see that their sum is precisely the LHS of (4).
A similar argument holds for even $h$. ∎
@@ IV-B General Lower Bounds
In particular, each interval has size $h$ and for all $\ell\in[h-1]$, we have that $I_{\ell}\cap I_{\ell+1}=\{\ell(h-2)+1\}$. Furthermore, let $J_{1},\ldots,J_{h(h-1)}\subseteq[n]$ be a partition of $[n]$ such that each set has size at least $N:=\lfloor\frac{n}{h(h-1)}\rfloor$.
@@
For any distinct $C_{\ell,j_{1},\ldots,j_{h-1}},C_{\ell,j^{\prime}_{1},\ldots,j^{\prime}_{h-1}}\in\mathcal{F}_{\ell}$, we have that the vectors $\Sigma_{H}(C_{\ell,j_{1},\ldots,j_{h-1}})$ and $\Sigma_{H}(C_{\ell,j^{\prime}_{1},\ldots,j^{\prime}_{h-1}})\}$ are linearly independent (i.e., projectively distinct).
@@
We can thus, define $A_{\ell}:=\{\Sigma_{H}(C):C\in\mathcal{F}_{\ell}\}$ for all $\ell\in[h]$. By Claim 22, each $A_{\ell}$ has size at least $N^{h}$ and its elements are projectively distinct. Furthermore, by Claim 23 any choice of $a_{1}\in A_{1},\ldots,a_{h}\in A_{h}$ are linearly independent, and thus $A_{1},\ldots,A_{h}$ are disjoint. We can apply Proposition 21 to get $q\geq\frac{N^{h-1}}{h-1}-4=\Omega(n^{h-1}/h^{2h-1})$, as desired. ∎
@@ IV-C Near-monotonicity in $h$
For any $m,n,h\geq h^{\prime}\geq 1$, we have that
@@
We claim that $H^{\prime}$ is MR. Consider any $h^{\prime}$ cycles $C^{\prime}_{1},\ldots,C^{\prime}_{h^{\prime}}\subseteq[m^{\prime}]\times[n^{\prime}]$ for which there is a spanning tree $T^{\prime}\subseteq[m^{\prime}]\times[n^{\prime}]$ such that $C^{\prime}_{\ell^{\prime}}\setminus T^{\prime}$ is a unique singleton for all $\ell^{\prime}\in[h^{\prime}]$. By Proposition 5, that $\Sigma_{H^{\prime}}(C^{\prime}_{1}),\ldots,\Sigma_{H^{\prime}}(C^{\prime}_{h^{\prime}})$ are linearly independent. However, by definition of $d^{i,j}$, we know that
@@ Acknowledgments
We thank Venkatesan Guruswami and Sergey Yekhanin for valuable comments on this manuscript.

### v2

@@ IV-A $n^{\lceil h/2\rceil}$ Lower Bound for the Gabidulin Construction
For the lower bound, we assume without loss of generality that $m=2$. Label the two left vertices as $\alpha$ and $\beta$ and label the right vertices as $1,\ldots,n$. If $h$ is odd then we claim that for any $(h+1)/2$ vertices $j_{1},\ldots,j_{(h+1)/2}\in[n]$ the sum of $\gamma(\alpha,j_{1})-\gamma(\beta,j_{1})+\cdots+\gamma(\alpha,j_{(h+1)/2})-\gamma(\beta,j_{(h+1)/2})$ is distinct for distinct choices of right vertices. We note this claim would then force the field size to be at least $\binom{n}{(h+1)/2}$.
@@
$\gamma(\alpha,i_{1})-\gamma(\beta,i_{1})+\ldots+\gamma(\alpha,i_{k})-\gamma(\beta,i_{k})-(\gamma(\alpha,j_{1})-\gamma(\beta,j_{1})+\ldots+\gamma(\alpha,j_{k})-\gamma(\beta,j_{k}))=0$ (4)
for $k\leq(h+1)/2$. Let $T$ be the following spanning tree of the subgraph induced by the aforementioned edges:
$T=\{(\alpha,i_{1}),\ldots,(\alpha,i_{k}),(\alpha,j_{1}),\ldots,(\alpha,j_{k}),(\beta,i_{1})\}.$
For any vertex $x\in\{i_{2},\ldots,i_{k},j_{1},\ldots,j_{k}\}$, the edge $(\beta,x)$ forms with $T$ a unique cycle
$C_{x}:=\{(\alpha,i_{1}),(\beta,i_{1}),(\beta,x),(\alpha,x)\}.$
In particular (after making an arbitrary sign choice), we have that
$\gamma(C_{x})=\gamma(\alpha,i_{1})-\gamma(\beta,i_{1})+\gamma(\beta,x)-\gamma(\alpha,x).$
Since our Gabidulin construction is MR (n,m,a=1,b=1,h=2k1)(n,m,a=1,b=1,h=2k-1), we have that $\{\gamma(C_{x}):x\in\{i_{2},\ldots,i_{k},j_{1},\ldots,j_{k}\}\}$ are linearly independent. In particular,
$0$ $\neq\sum_{x\in\{j_{1},\ldots,j_{k}\}}\gamma(C_{x})-\sum_{x\in\{i_{2},\ldots,i_{k}\}}\gamma(C_{x})$
$=\gamma(\alpha,i_{1})-\gamma(\beta,i_{1})+\sum_{x\in\{i_{2},\ldots,i_{k}\}}(\gamma(\alpha,x)-\gamma(\beta,x))-\sum_{x\in\{j_{1},\ldots,j_{k}\}}(\gamma(\alpha,x)-\gamma(\beta,x))$
[…]
a contradiction.
A similar argument holds for even $h$ by picking $k=h/2$ and noting that $h\geq 2k-1$. ∎
@@ IV-B General Lower Bounds
In particular, each interval has size h1h-1 and for all $\ell\in[h-1]$, we have that $I_{\ell}\cap I_{\ell+1}=\{\ell(h-2)+1\}$. Furthermore, let $J_{1},\ldots,J_{h(h-1)}\subseteq[n]$ be a partition of $[n]$ such that each set has size at least $N:=\lfloor\frac{n}{h(h-1)}\rfloor$.
@@
For any distinct $C_{\ell,j_{1},\ldots,j_{h-1}},C_{\ell,j^{\prime}_{1},\ldots,j^{\prime}_{h-1}}\in\mathcal{F}_{\ell}$, we have that the vectors $\Sigma_{H}(C_{\ell,j_{1},\ldots,j_{h-1}})$ and $\Sigma_{H}(C_{\ell,j^{\prime}_{1},\ldots,j^{\prime}_{h-1}})$ are linearly independent (i.e., projectively distinct).
@@
We can thus, define $A_{\ell}:=\{\Sigma_{H}(C):C\in\mathcal{F}_{\ell}\}$ for all $\ell\in[h]$. By Claim 24, each $A_{\ell}$ has size at least $N^{h-1}$ and its elements are projectively distinct. Furthermore, by Claim 25 any choice of $a_{1}\in A_{1},\ldots,a_{h}\in A_{h}$ are linearly independent, and thus $A_{1},\ldots,A_{h}$ are disjoint. We can apply Proposition 23 to get $q\geq\frac{N^{h-1}}{h-1}-4=\Omega(n^{h-1}/h^{2h-1})$, as desired. ∎
@@ IV-C Near-monotonicity in $h$
A general question one may ask about $q(m,n,a,b,h)$ is the dependence on the field size as a function of the number of global parity checks $h$. Intuitively, demanding more global parity checks appears to be a stronger property of the code and thus the field size should grow as a function of $h$. However, rather surprisingly, no such result is known. The primary issue is that if $h>h^{\prime}$, an MR grid code with parameters $(m,n,a,b,h)$ does not immediately imply an MR grid code with parameters $(m,n,a,b,h^{\prime})$ as the former code has smaller dimension than the latter, and it is not clear how to make up for the deficit in dimension.
@@
For any $m\geq 3$, $h\geq h^{\prime}\geq 1$, and $n\geq h-h^{\prime}+2$, we have that
@@
We claim that $H^{\prime}$ is MR. Consider any $h^{\prime}$ cycles $C^{\prime}_{1},\ldots,C^{\prime}_{h^{\prime}}\subseteq[m^{\prime}]\times[n^{\prime}]$ for which there is a spanning tree $T^{\prime}\subseteq[m^{\prime}]\times[n^{\prime}]$ such that $C^{\prime}_{\ell^{\prime}}\setminus T^{\prime}$ has a unique single edge for all $\ell^{\prime}\in[h^{\prime}]$. By Proposition 7, it suffices to show that $\Sigma_{H^{\prime}}(C^{\prime}_{1}),\ldots,\Sigma_{H^{\prime}}(C^{\prime}_{h^{\prime}})$ are linearly independent. By definition of $d^{i,j}$, we know that
@@ Acknowledgments
We thank Venkatesan Guruswami and Sergey Yekhanin for valuable comments on this manuscript. We also thank anonymous reviewers for numerous suggestions improving the quality and correctness of this manuscript.